The Ksp value for Gd(OH)3 is 2.8*10^-23, the PH at which Gd(OH)3 begins to precipitate is:-
a.5.08    b.8.47     c.4.08

Grade:

1 Answers

Arun
5 years ago

Solution:

1) Write the dissociation equation:
Gd(OH)3 ⇆ Gd3+ + 3OH¯

2) Write the Ksp expression:
Ksp = [Gd3+] [OH¯]3

3) Let the concentration of each species is x.
Plug into the Ksp expression:
2.8 x 10¯23 = (x) (3x)3
                 = (x) (27x3)
​                 = 27 x4
x4 = 2.8 x 10-23/27
    = 0.1 x 10-23
    = 10-24
x = 10-6
 

4) Solve for [OH¯]:
[Gd3+] = x = 10¯6 
[OH¯] = 3x = 3 x 10¯6 M

5) Calculate the pH given the [OH¯]:
pOH = -log 3 x 10¯6 = 5.5
pH = 14 - pOH = 14 - 5.5 = 8.5

So, Gd(OH)3 will precipitate if the pH is 8.5 or higher.